# Logarithmic regret in the ergodic Avellaneda–Stoikov market making model

### v1

@@ A.13. Proof of Proposition 21
First, let $v_{r}^{\psi^{\kappa}}(q;\kappa^{\star})$ be the expected reward in the discounted infinite-time-horizon setting, where the true price sensitivity parameter is $\kappa^{\star}$ but the market maker uses the strategy $\psi^{\kappa}$ given by (18) with parameter $\kappa$, i.e.
(42) $v_{r}^{\psi^{\kappa}}(q;\kappa^{\star})=\mathbb{E}_{q}\Big{[}\int_{0}^{+\infty}e^{-rt}f(Q_{t}^{\psi^{\kappa};\kappa^{\star}},\psi^{\kappa};\kappa^{\star})\,\mathrm{d}t\Big{]},$
where $f(\cdot;\kappa^{\star})$ is given by (21). Then we claim that $v_{r}^{\psi^{\kappa}}(q;\kappa^{\star})$ satisfies the following linear system
(43) $0=-rv_{r}^{\psi^{\kappa}}(q;\kappa^{\star})+H\Big{(}q,\psi^{\kappa},(v_{r}^{\psi^{\kappa}}(q^{\prime};\kappa^{\star})-v_{r}^{\psi^{\kappa}}(q;\kappa^{\star}))_{q^{\prime}\in\{q-1,q+1\}};\kappa^{\star}\Big{)},\quad\forall q\in\Omega^{Q}\,.$
We would like to provide a sketch of proof for the claim. First, the existence of $v_{r}^{\psi^{\kappa}}(q;\kappa^{\star})$ defined by (42) can follow the proof in Section A.2 for Theorem 3 but substituting the Hamiltonian function $H$ (10) for $H(\cdot;\kappa^{\star})$ (27). Let $w(q)$ be a solution for the linear system (43). Consider
$X(s)=e^{-rs}w(Q_{s})+\int_{0}^{s}e^{-rt}f(Q_{t},\psi^{\kappa};\kappa^{\star})\,\mathrm{d}t\,.$
@@
$dX(s)=$ $\,e^{-rs}\Big{\{}-rw(Q_{s})+\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,+}_{s}}\left(w(Q_{s}-1)-w(Q_{s})\right)\bm{1}_{Q_{s}>\underline{q}}$
$\quad+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,-}_{s}}\left(w(Q_{s}+1)-w(Q_{s})\right)\bm{1}_{Q_{s}<\bar{q}}+f(Q_{s},\psi^{\kappa};\kappa^{\star})\Big{\}}ds$
$+e^{-rs}\Big{\{}\left(w(Q_{s}-1)-w(Q_{s})\right)\bm{1}_{Q_{s}>\underline{q}}d\tilde{N}_{s}^{+}+\left(w(Q_{s}+1)-w(Q_{s})\right)\bm{1}_{Q_{s}<\bar{q}}d\tilde{N}_{s}^{-}\Big{\}}$
@@
$+e^{-rs}\Big{\{}\left(w(Q_{s}-1)-w(Q_{s})\right)\bm{1}_{Q_{s}>\underline{q}}d\tilde{N}_{s}^{+}+\left(w(Q_{s}+1)-w(Q_{s})\right)\bm{1}_{Q_{s}<\bar{q}}d\tilde{N}_{s}^{-}\Big{\}}.$
where the last equality comes from the definition of $H(\cdot;\kappa^{\star})$ (27) and $f(\cdot;\kappa^{\star})$ (21). Take the integral and expectation for $X(s)$, we have
@@
$\begin{split}w(q)=\mathbb{E}\big{[}X(0)\big{|}Q_{0}=q\big{]}&=\mathbb{E}\big{[}X(T)\big{|}Q_{0}=q\big{]}\\ &=\mathbb{E}_{q}\Big{[}\int_{0}^{T}e^{-rt}f(Q_{t},\psi^{\kappa};\kappa^{\star})\,\mathrm{d}t+e^{-rT}w(Q_{T})\Big{]}.\end{split}$
@@
$w(q)=\mathbb{E}_{q}\Big{[}\int_{0}^{+\infty}e^{-rt}f(Q_{t},\psi^{\kappa};\kappa^{\star})\,\mathrm{d}t\Big{]}.$
@@
Now, we would like to to show that given $\gamma(\kappa;\kappa^{\star})$ defined by (30), it holds that
(44) $\lim_{r\rightarrow 0}rv_{r}^{\psi^{\kappa}}(q;\kappa^{\star})=\gamma(\kappa;\kappa^{\star}),\quad\forall q\in\Omega^{Q}\,,$
where $v_{r}^{\psi^{\kappa}}(q;\kappa^{\star})$ is defined by (42).
@@
    $\exists\,C_{1}\in\mathbb{R}^{+}$ such that $\left|rv_{r}^{\psi^{\kappa}}(q;\kappa^{\star})\right|\leq C_{1}$ for $\forall q\in\Omega^{Q}$ and $r\in\mathbb{R}^{+}$.
@@
    $\exists\,C_{2}\in\mathbb{R}^{+}$ such that $\left|v_{r}^{\psi^{\kappa}}(\hat{q};\kappa^{\star})-v_{r}^{\psi^{\kappa}}(q;\kappa^{\star})\right|\leq C_{2}|\hat{q}-q|$ for $\forall q,\hat{q}\in\Omega^{Q}$ and $r\in\mathbb{R}^{+}$.
As $\psi^{\kappa}\subset\mathcal{A}$, i.e. the collection of all Markov controls optimal for the ergodic control problem is a subset of the admissible control, the running reward function $f(\cdot;\kappa^{\star})$ is bounded. By (21), we have
(45) $\begin{split}f(Q_{t},\psi^{\kappa};\kappa^{\star})&=\lambda^{+}\psi^{\kappa,+}e^{-\kappa^{\star}\psi^{\kappa,+}}+\lambda^{-}\psi^{\kappa,-}e^{-\kappa^{\star}\psi^{\kappa,-}}-\phi(Q_{t})^{2}\\ &\leq\sup_{\psi^{\kappa}\in\overline{\mathbb{R}}^{2}}\big{(}\lambda^{+}\psi^{\kappa,+}e^{-\kappa^{\star}\psi^{\kappa,+}}+\lambda^{-}\psi^{\kappa,-}e^{-\kappa^{\star}\psi^{\kappa,-}}\big{)}\\ &=\overline{C}\,\end{split}$
@@
(46) $\begin{split}f(Q_{t},\psi^{\kappa};\kappa^{\star})&=\lambda^{+}\psi^{\kappa,+}e^{-\kappa^{\star}\psi^{\kappa,+}}+\lambda^{-}\psi^{\kappa,-}e^{-\kappa^{\star}\psi^{\kappa,-}}-\phi(Q_{t})^{2}\\ &\geq\inf_{\psi^{\kappa}}\big{(}\lambda^{+}\psi^{\kappa,+}e^{-\kappa^{\star}\psi^{\kappa,+}}+\lambda^{-}\psi^{\kappa,-}e^{-\kappa^{\star}\psi^{\kappa,-}}\big{)}-\phi\max(\bar{q},\underline{q})^{2}\\ &=\underline{C}\,.\end{split}$
@@
$\begin{split}v_{r}^{\psi^{\kappa}}(q;\kappa^{\star})&=\mathbb{E}_{q}\Big{[}\int_{0}^{+\infty}e^{-rt}f(Q_{t}^{\psi^{\kappa};\kappa^{\star}},\psi^{\kappa};\kappa^{\star})\,\mathrm{d}t\Big{]}\geq\mathbb{E}_{q}\Big{[}\int_{0}^{+\infty}e^{-rt}\underline{C}\,\mathrm{d}t\Big{]}=\frac{\underline{C}}{r}\,.\end{split}$
@@
$\begin{split}v_{r}^{\psi^{\kappa}}(q;\kappa^{\star})\leq\mathbb{E}_{q}\Big{[}\int_{0}^{+\infty}e^{-rt}\overline{C}\,\mathrm{d}t\Big{]}=\frac{\bar{C}}{r}\,.\end{split}$
@@
To prove Lemma 37 (2), we define the stopping time $\tau(q,\hat{q})$ for $Q_{t}^{\psi^{\kappa};\kappa^{\star}}$ with $Q_{0}=q$ as
$\tau:=\inf\{t\left.\right|Q_{t}^{\psi^{\kappa};\kappa^{\star}}=\hat{q}\}\,.$
With the fact that, $\forall q,\hat{q}\in\Omega^{Q}$, the transition probability for $\hat{q}$ from $q$, i.e. $\bm{P}(q,\hat{q})>0$, see the more detailed discussion in Section A.10, it implies that $\mathbb{E}[Q_{\tau}]<+\infty$.
By (45) and (46), there exists $\overline{C},\underline{C}$ such that $0\leq f(q;\psi^{\kappa};\kappa^{\star})-\underline{C}\leq\overline{C}-\underline{C}$ for any $q\in\Omega^{Q}$. Therefore,
$\begin{split}v_{r}^{\psi^{\kappa}}(q;\kappa^{\star})-\frac{\underline{C}}{r}&=\mathbb{E}_{q}\Big{[}\int_{0}^{\tau}e^{-rt}\big{(}f(Q^{\psi^{\kappa};\kappa^{\star}}_{t},\psi^{\kappa};\kappa^{\star})-\underline{C}\big{)}\,\mathrm{d}t+e^{-r\tau}\big{(}v_{r}^{\psi^{\kappa}}(\hat{q};\kappa^{\star})-\frac{\underline{C}}{r}\big{)}\Big{]}\\ &\leq\mathbb{E}\Big{[}\int_{0}^{\tau}e^{-rt}\big{(}\overline{C}-\underline{C}\big{)}\,\mathrm{d}t\Big{]}+\mathbb{E}\big{[}e^{-r\tau}\big{]}\big{(}v^{\psi^{\kappa};\kappa^{\star}}_{r}(\hat{q})-\frac{\underline{C}}{r}\big{)}\\ &\leq(\bar{C}-\underline{C})\mathbb{E}[\tau]+v^{\psi^{\kappa};\kappa^{\star}}_{r}(\hat{q})-\frac{\underline{C}}{r}\,.\end{split}$
@@
$\begin{split}\frac{v^{\psi^{\kappa};\kappa^{\star}}_{r}(q)-v^{\psi^{\kappa};\kappa^{\star}}_{r}(\hat{q})}{|q-\hat{q}|}&\leq\frac{1}{|q-\hat{q}|}\Big{(}(\bar{C}-\underline{C})\mathbb{E}[\tau]\Big{)}\leq(\bar{C}-\underline{C})\mathbb{E}[\tau],\,\forall q,\hat{q}\in\Omega^{Q},q\neq\hat{q}.\end{split}$
Since $\mathbb{E}[\tau]<+\infty$, we conclude that $\frac{v^{\psi^{\kappa};\kappa^{\star}}_{r}(q)-v^{\psi^{\kappa};\kappa^{\star}}_{r}(\hat{q})}{|q-\hat{q}|}$ is bounded from above. By simply changing the order of $q$ and $\hat{q}$, we conclude the lower boundedness. Hence, we can find $C_{2}\in\mathbb{R}^{+}$ such that
$\left|v^{\psi^{\kappa};\kappa^{\star}}_{r}(q)-v^{\psi^{\kappa};\kappa^{\star}}_{r}(\hat{q})\right|\leq C_{2}|\hat{q}-q|,\,\forall q,\hat{q}\in\Omega^{Q},r\in\mathbb{R}^{+}\,.$
With the fact that $\Big{|}rv_{r}^{\psi^{\kappa}}(q;\kappa^{\star})\Big{|}$ and $\Big{|}v_{r}^{\psi^{\kappa}}(\hat{q};\kappa^{\star})-v_{r}^{\psi^{\kappa}}(q;\kappa^{\star})\Big{|}$ are bounded, we can follow the discussion in Appendix A.3, i.e. consider a sequence $(r_{n})_{n\in\mathbb{N}}$ converging towards 00 such that the sequences $\big{(}r_{n}v_{r_{n}}^{\psi^{\kappa};\kappa^{\star}}(q)\big{)}_{n\in\mathbb{N}}$ and $\big{(}v_{r_{n}}^{\psi^{\kappa};\kappa^{\star}}(q)-v_{r_{n}}^{\psi^{\kappa};\kappa^{\star}}(0)\big{)}_{n\in\mathbb{N}}$ are convergent for $\forall q\in\Omega^{Q}$, then show that there exists $\gamma(\kappa;\kappa^{\star})\in\mathbb{R}$ such that $\gamma(\kappa;\kappa^{\star})=\lim_{r\rightarrow 0}rv_{r}^{\psi^{\kappa}}(q;\kappa^{\star})$. Again by substituting $H$ (10) for $H(\cdot;\kappa^{\star})$ in the proof of Theorem 5 in Appendix A.4, we can easily conclude that $\gamma(\kappa;\kappa^{\star})$ given by (44) also satisfies
$\gamma(\kappa;\kappa^{\star})=\lim_{T\rightarrow+\infty}\frac{1}{T}v_{0}^{\psi^{\kappa}}(t=0,q;T;\kappa^{\star})\,.$
Let us define $\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star})=\lim_{r\rightarrow 0}\big{(}v_{r}^{\psi^{\kappa}}(q;\kappa^{\star})-v_{r}^{\psi^{\kappa}}(0;\kappa^{\star})\big{)}$ for $q\in\Omega^{Q}$. By the convergent of $\big{(}v_{r_{n}}^{\psi^{\kappa};\kappa^{\star}}(q)-v_{r_{n}}^{\psi^{\kappa};\kappa^{\star}}(0)\big{)}_{n\in\mathbb{N}}$ under the sequence $(r_{n})_{n\in\mathbb{N}}$ converging towards 00, we know that $\hat{v}^{\psi^{\kappa}}(q;\kappa^{\star})$ is well defined.
@@
$0=-\gamma(\kappa;\kappa^{\star})+H\Big{(}q,\psi^{\kappa},(\hat{v}^{\psi^{\kappa},\kappa^{\star}}(q^{\prime})-\hat{v}^{\psi^{\kappa},\kappa^{\star}}(q))_{q^{\prime}\in\{q-1,q+1\}};\kappa^{\star}\Big{)}\,,\,\forall q\in\Omega^{Q}.$
A.14. Proof of Proposition 22

### v2

@@ A.12. Proof of Proposition 19
First, let $v_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})$ be the expected reward in the discounted infinite-time-horizon setting, where the true price sensitivity parameter is $\kappa^{\ast}$ but the market maker uses the strategy $\psi^{\kappa}$ given by (19) with parameter $\kappa$, i.e.
(45) $v_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})=\mathbb{E}_{q}\Big{[}\int_{0}^{+\infty}e^{-rt}f(Q_{t}^{\psi^{\kappa};\kappa^{\ast}},\psi^{\kappa};\kappa^{\ast})\,\mathrm{d}t\Big{]},$
where $f(\cdot;\kappa^{\ast})$ is given by (22). Then we claim that $v_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})$ satisfies the following linear system
(46) $0=-rv_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})+H\Big{(}q,\psi^{\kappa},(v_{r}^{\psi^{\kappa}}(q^{\prime};\kappa^{\ast})-v_{r}^{\psi^{\kappa}}(q;\kappa^{\ast}))_{q^{\prime}\in\{q-1,q+1\}};\kappa^{\ast}\Big{)},\quad\forall q\in\Omega^{Q}\,.$
We would like to provide a sketch of proof for the claim. First, the existence of $v_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})$ defined by (45) can follow the proof in Section A.2 for Theorem 3 but substituting the Hamiltonian function $H$ (10) for $H(\cdot;\kappa^{\ast})$ (31). Let $w(q)$ be a solution for the linear system (46). Consider
$X(s)=e^{-rs}w(Q_{s})+\int_{0}^{s}e^{-rt}f(Q_{t},\psi^{\kappa};\kappa^{\ast})\,\mathrm{d}t\,.$
@@
$dX(s)=$ $\,e^{-rs}\Big{\{}-rw(Q_{s})+\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}_{s}}\left(w(Q_{s}-1)-w(Q_{s})\right)\mathds{1}_{Q_{s}>\underline{q}}$
$\quad+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}_{s}}\left(w(Q_{s}+1)-w(Q_{s})\right)\mathds{1}_{Q_{s}<\bar{q}}+f(Q_{s},\psi^{\kappa};\kappa^{\ast})\Big{\}}ds$
$+e^{-rs}\Big{\{}\left(w(Q_{s}-1)-w(Q_{s})\right)\mathds{1}_{Q_{s}>\underline{q}}d\tilde{N}_{s}^{+}+\left(w(Q_{s}+1)-w(Q_{s})\right)\mathds{1}_{Q_{s}<\bar{q}}d\tilde{N}_{s}^{-}\Big{\}}$
@@
$+e^{-rs}\Big{\{}\left(w(Q_{s}-1)-w(Q_{s})\right)\mathds{1}_{Q_{s}>\underline{q}}d\tilde{N}_{s}^{+}+\left(w(Q_{s}+1)-w(Q_{s})\right)\mathds{1}_{Q_{s}<\bar{q}}d\tilde{N}_{s}^{-}\Big{\}}.$
where the last equality comes from the definition of $H(\cdot;\kappa^{\ast})$ (31) and $f(\cdot;\kappa^{\ast})$ (22). Take the integral and expectation for $X(s)$, we have
@@
$\begin{split}w(q)=\mathbb{E}\big{[}X(0)\big{|}Q_{0}=q\big{]}&=\mathbb{E}\big{[}X(T)\big{|}Q_{0}=q\big{]}\\ &=\mathbb{E}_{q}\Big{[}\int_{0}^{T}e^{-rt}f(Q_{t},\psi^{\kappa};\kappa^{\ast})\,\mathrm{d}t+e^{-rT}w(Q_{T})\Big{]}.\end{split}$
@@
$w(q)=\mathbb{E}_{q}\Big{[}\int_{0}^{+\infty}e^{-rt}f(Q_{t},\psi^{\kappa};\kappa^{\ast})\,\mathrm{d}t\Big{]}.$
@@
Now, we would like to to show that given $\gamma(\kappa;\kappa^{\ast})$ defined by (34), it holds that
(47) $\lim_{r\rightarrow 0}rv_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})=\gamma(\kappa;\kappa^{\ast}),\quad\forall q\in\Omega^{Q}\,,$
where $v_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})$ is defined by (45).
@@
    $\exists\,C_{1}\in\mathbb{R}^{+}$ such that $\left|rv_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})\right|\leq C_{1}$ for $\forall q\in\Omega^{Q}$ and $r\in\mathbb{R}^{+}$.
@@
    $\exists\,C_{2}\in\mathbb{R}^{+}$ such that $\left|v_{r}^{\psi^{\kappa}}(\hat{q};\kappa^{\ast})-v_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})\right|\leq C_{2}|\hat{q}-q|$ for $\forall q,\hat{q}\in\Omega^{Q}$ and $r\in\mathbb{R}^{+}$.
As $\psi^{\kappa}\subset\mathcal{A}$, i.e. the collection of all Markov controls optimal for the ergodic control problem is a subset of the admissible control, the running reward function $f(\cdot;\kappa^{\ast})$ is bounded. By (22), we have
(48) $\begin{split}f(Q_{t},\psi^{\kappa};\kappa^{\ast})&=\lambda^{+}\psi^{\kappa,+}e^{-\kappa^{\ast}\psi^{\kappa,+}}+\lambda^{-}\psi^{\kappa,-}e^{-\kappa^{\ast}\psi^{\kappa,-}}-\phi(Q_{t})^{2}\\ &\leq\sup_{\psi^{\kappa}\in\overline{\mathbb{R}}^{2}}\big{(}\lambda^{+}\psi^{\kappa,+}e^{-\kappa^{\ast}\psi^{\kappa,+}}+\lambda^{-}\psi^{\kappa,-}e^{-\kappa^{\ast}\psi^{\kappa,-}}\big{)}\\ &=\overline{C}\,\end{split}$
@@
(49) $\begin{split}f(Q_{t},\psi^{\kappa};\kappa^{\ast})&=\lambda^{+}\psi^{\kappa,+}e^{-\kappa^{\ast}\psi^{\kappa,+}}+\lambda^{-}\psi^{\kappa,-}e^{-\kappa^{\ast}\psi^{\kappa,-}}-\phi(Q_{t})^{2}\\ &\geq\inf_{\psi^{\kappa}}\big{(}\lambda^{+}\psi^{\kappa,+}e^{-\kappa^{\ast}\psi^{\kappa,+}}+\lambda^{-}\psi^{\kappa,-}e^{-\kappa^{\ast}\psi^{\kappa,-}}\big{)}-\phi\max(\bar{q},\underline{q})^{2}\\ &=\underline{C}\,.\end{split}$
@@
$\begin{split}v_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})&=\mathbb{E}_{q}\Big{[}\int_{0}^{+\infty}e^{-rt}f(Q_{t}^{\psi^{\kappa};\kappa^{\ast}},\psi^{\kappa};\kappa^{\ast})\,\mathrm{d}t\Big{]}\geq\mathbb{E}_{q}\Big{[}\int_{0}^{+\infty}e^{-rt}\underline{C}\,\mathrm{d}t\Big{]}=\frac{\underline{C}}{r}\,.\end{split}$
@@
$\begin{split}v_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})\leq\mathbb{E}_{q}\Big{[}\int_{0}^{+\infty}e^{-rt}\overline{C}\,\mathrm{d}t\Big{]}=\frac{\bar{C}}{r}\,.\end{split}$
@@
To prove Lemma 39 (2), we define the stopping time $\tau(q,\hat{q})$ for the process $Q_{t}^{\psi^{\kappa};\kappa^{\ast}}$ with initial condition $Q_{0}=q$ as
$\tau:=\inf\{t\left.\right|Q_{t}^{\psi^{\kappa};\kappa^{\ast}}=\hat{q}\}\,.$
Since the dynamics of $Q_{t}^{\psi^{\kappa};\kappa^{\ast}}$ can be equivalently represented by a continuous-time Markov chain that is irreducible and recurrent (see the detailed discussion in Section A.15), it follows that $\mathbb{E}[\tau]<+\infty$.
By (48) and (49), there exists $\overline{C},\underline{C}$ such that $0\leq f(q;\psi^{\kappa};\kappa^{\ast})-\underline{C}\leq\overline{C}-\underline{C}$ for any $q\in\Omega^{Q}$. Therefore,
$\begin{split}v_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})-\frac{\underline{C}}{r}&=\mathbb{E}_{q}\Big{[}\int_{0}^{\tau}e^{-rt}\big{(}f(Q^{\psi^{\kappa};\kappa^{\ast}}_{t},\psi^{\kappa};\kappa^{\ast})-\underline{C}\big{)}\,\mathrm{d}t+e^{-r\tau}\big{(}v_{r}^{\psi^{\kappa}}(\hat{q};\kappa^{\ast})-\frac{\underline{C}}{r}\big{)}\Big{]}\\ &\leq\mathbb{E}\Big{[}\int_{0}^{\tau}e^{-rt}\big{(}\overline{C}-\underline{C}\big{)}\,\mathrm{d}t\Big{]}+\mathbb{E}\big{[}e^{-r\tau}\big{]}\big{(}v^{\psi^{\kappa};\kappa^{\ast}}_{r}(\hat{q})-\frac{\underline{C}}{r}\big{)}\\ &\leq(\bar{C}-\underline{C})\mathbb{E}[\tau]+v^{\psi^{\kappa};\kappa^{\ast}}_{r}(\hat{q})-\frac{\underline{C}}{r}\,.\end{split}$
@@
$\begin{split}\frac{v^{\psi^{\kappa};\kappa^{\ast}}_{r}(q)-v^{\psi^{\kappa};\kappa^{\ast}}_{r}(\hat{q})}{|q-\hat{q}|}&\leq\frac{1}{|q-\hat{q}|}\Big{(}(\bar{C}-\underline{C})\mathbb{E}[\tau]\Big{)}\leq(\bar{C}-\underline{C})\mathbb{E}[\tau],\,\forall q,\hat{q}\in\Omega^{Q},q\neq\hat{q}.\end{split}$
Since $\mathbb{E}[\tau]<+\infty$, we conclude that $\frac{v^{\psi^{\kappa};\kappa^{\ast}}_{r}(q)-v^{\psi^{\kappa};\kappa^{\ast}}_{r}(\hat{q})}{|q-\hat{q}|}$ is bounded from above. By simply changing the order of $q$ and $\hat{q}$, we conclude the lower boundedness. Hence, we can find $C_{2}\in\mathbb{R}^{+}$ such that
$\left|v^{\psi^{\kappa};\kappa^{\ast}}_{r}(q)-v^{\psi^{\kappa};\kappa^{\ast}}_{r}(\hat{q})\right|\leq C_{2}|\hat{q}-q|,\,\forall q,\hat{q}\in\Omega^{Q},r\in\mathbb{R}^{+}\,.$
With the fact that $\Big{|}rv_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})\Big{|}$ and $\Big{|}v_{r}^{\psi^{\kappa}}(\hat{q};\kappa^{\ast})-v_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})\Big{|}$ are bounded, we can follow the discussion in Appendix A.3, i.e. consider a sequence $(r_{n})_{n\in\mathbb{N}}$ converging towards 00 such that the sequences $\big{(}r_{n}v_{r_{n}}^{\psi^{\kappa};\kappa^{\ast}}(q)\big{)}_{n\in\mathbb{N}}$ and $\big{(}v_{r_{n}}^{\psi^{\kappa};\kappa^{\ast}}(q)-v_{r_{n}}^{\psi^{\kappa};\kappa^{\ast}}(0)\big{)}_{n\in\mathbb{N}}$ are convergent for $\forall q\in\Omega^{Q}$, then show that there exists $\gamma(\kappa;\kappa^{\ast})\in\mathbb{R}$ such that $\gamma(\kappa;\kappa^{\ast})=\lim_{r\rightarrow 0}rv_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})$. Again by substituting $H$ (10) for $H(\cdot;\kappa^{\ast})$ in the proof of Theorem 5 in Appendix A.4, we can easily conclude that $\gamma(\kappa;\kappa^{\ast})$ given by (47) also satisfies
$\gamma(\kappa;\kappa^{\ast})=\lim_{T\rightarrow+\infty}\frac{1}{T}v_{0}^{\psi^{\kappa}}(t=0,q;T;\kappa^{\ast})\,.$
Let us define $\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast})=\lim_{r\rightarrow 0}\big{(}v_{r}^{\psi^{\kappa}}(q;\kappa^{\ast})-v_{r}^{\psi^{\kappa}}(0;\kappa^{\ast})\big{)}$ for $q\in\Omega^{Q}$. By the convergent of $\big{(}v_{r_{n}}^{\psi^{\kappa};\kappa^{\ast}}(q)-v_{r_{n}}^{\psi^{\kappa};\kappa^{\ast}}(0)\big{)}_{n\in\mathbb{N}}$ under the sequence $(r_{n})_{n\in\mathbb{N}}$ converging towards 00, we know that $\hat{v}^{\psi^{\kappa}}(q;\kappa^{\ast})$ is well defined.
@@
$0=-\gamma(\kappa;\kappa^{\ast})+H\Big{(}q,\psi^{\kappa},(\hat{v}^{\psi^{\kappa},\kappa^{\ast}}(q^{\prime})-\hat{v}^{\psi^{\kappa},\kappa^{\ast}}(q))_{q^{\prime}\in\{q-1,q+1\}};\kappa^{\ast}\Big{)}\,,\,\forall q\in\Omega^{Q}.$
A.13. Proof of Proposition 20